# AdaptGCD: Multi-Expert Adapter Tuning for Generalized Category Discovery

### v1

@@ Abstract
Different from the traditional semi-supervised learning paradigm that is constrained by the close-world assumption, Generalized Category Discovery (GCD) presumes that the unlabeled dataset contains new categories not appearing in the labeled set, and aims to not only classify old categories but also discover new categories in the unlabeled data. Existing studies on GCD typically devote to transferring the general knowledge from the self-supervised pretrained model to the target GCD task via some fine-tuning strategies, such as partial tuning and prompt learning. Nevertheless, these fine-tuning methods fail to make a sound balance between the generalization capacity of pretrained backbone and the adaptability to the GCD task. To fill this gap, in this paper, we propose a novel adapter-tuning-based method named AdaptGCD, which is the first work to introduce the adapter tuning into the GCD task and provides some key insights expected to enlighten future research. Furthermore, considering the discrepancy of supervision information between the old and new classes, a multi-expert adapter structure equipped with a route assignment constraint is elaborately devised, such that the data from old and new classes are separated into different expert groups. Extensive experiments are conducted on 7 widely-used datasets. The remarkable improvements in performance highlight the effectiveness of our proposals.
1 Introduction
@@
While recent semi-supervised studies DBLP:conf/nips/SohnBCZZRCKL20 ; DBLP:conf/nips/BerthelotCGPOR19 ; DBLP:conf/nips/ZhangWHWWOS21 ; DBLP:conf/nips/LiWLYY0L23 ; DBLP:conf/nips/SunSL23 ; DBLP:conf/nips/Huang0Y0L23 show a substantial progression, they are typically constrained by the closed-world assumption, i.e., labels are provided for each category that the model needs to classify DBLP:conf/iclr/CaoBL22 . The constraint hinders the scalability of these methods. To relax this assumption, Generalized Category Discovery (GCD) DBLP:conf/cvpr/gcdVazeHVZ22 presumes that the unlabeled dataset contains new categories not appearing in the labeled set, as shown in Figure 1 (a). The goal of GCD is not only to classify old categories but also to discover new categories in the unlabeled data.
@@
The primary challenge in GCD is learning the expressive representation applicable to both old and new classes DBLP:conf/cvpr/gcdVazeHVZ22 . Due to the scarcity of labels in GCD tasks, the model trained from scratch is prone to overfit the labeled data, consequently disrupting the generality of representation wang2024sptnet . As an alternative, the existing studies in GCD broadly finetune the self-supervised pretrained models, like DINO DBLP:conf/iccv/CaronTMJMBJ21 , to transfer the strong general knowledge from the pretrained backbone to the target GCD data. In this process, a key concern is to strike a balance between the generalization capacity from the upstream knowledge and the adaptability to the downstream GCD task, which is largely influenced by the fine-tuning strategy DBLP:journals/corr/abs-2403-14608 ; DBLP:conf/nips/LiuTMMHBR22 ; DBLP:conf/eccv/vptJiaTCCBHL22 ; DBLP:conf/acl/LiL20 .
Figure 1: (a) The description of the generalized category discovery (GCD) task. (b) The framework of our proposed AdaptGCD. It contains two critical modules: the multi-expert adapter (MEA) structure and the route assignment constraint. The MEA introduces multiple adapter experts and the route assignment constraint guides the allocation of these experts.
@@
Up to now, two fine-tuning strategies are widely utilized in GCD: the partial tuning and the prompt tuning. The straightforward partial tuning strategy only updates the parameters from the last few blocks while the rest parameters in the backbone are frozen DBLP:conf/cvpr/gcdVazeHVZ22 . Such approaches potentially impair the generalization ability since the pretrained knowledge embodied in the backbone is disrupted. Conversely, prompt tuning methods DBLP:conf/eccv/vptJiaTCCBHL22 ; DBLP:conf/nips/WangC0H23 like PromptCAL DBLP:conf/cvpr/promptcalgcdZhangKSNCK23 and SPTNet wang2024sptnet freeze all network parameters in the backbone and incorporate a learnable prompt at the input for training. Although effectively preserving the knowledge of pretrained network, unfortunately, it is demonstrated that the prompt-based approaches are not robust to bad data quality wang2024sptnet , and at the same time present limited adaptability when the downstream task differs from the upstream pretraining task DBLP:conf/eccv/vptJiaTCCBHL22 , thus making their performance on GCD task far from satisfactory.
@@
Luckily, as one of the most characteristic fine-tuning approaches, the adapter tuning DBLP:conf/nips/adaptformerChenGTWSWL22 ; DBLP:conf/nips/RebuffiBV17 ; DBLP:conf/nips/LeiBBALZ0ZWLZC23 ; DBLP:conf/nips/Page-CacciaPSPR23 ; DBLP:conf/nips/ChengSXW0GS23 may provide a promising solution to the issues mentioned above. The adapter tuning leaves the pretrained model fixed and brings in a modest number of task-specific parameters via bottleneck adapter modules DBLP:journals/corr/gloraabs-2306-07967 . Obviously, on mechanism, the adapter naturally fits for the GCD task. It could preserve the pretrained knowledge nondestructively DBLP:conf/acl/WangTDWHJCJZ21 ; DBLP:journals/taslp/HouZWWQXS22 ; DBLP:conf/miccai/ZhangHZZZW23 , as well as provide flexibility for inserting domain knowledge of diverse downstream GCD tasks DBLP:conf/acl/Ye020 ; DBLP:conf/cvpr/LiLB22 ; DBLP:conf/aaai/XinDWLY24 . Surprisingly, the adapter has not yet attracted much attention on the GCD task. When introducing the adapter into GCD, an additional concern arises accordingly, i.e., the information imbalance between old and new classes within the adapter DBLP:conf/iclr/CaoBL22 . Due to the existence of supervised loss for labeled data, the adapter receives more supervision from old classes than the new ones, thereby resulting in a bias towards old classes. Under the adapter framework, how to render old classes and new classes non-interfering in the GCD task deserves further exploration.
@@
In this study, we propose a novel multi-expert adapter tuning method AdaptGCD for generalized category discovery as shown in Figure 1 (b). Unlike the previous partial tuning or prompt-like methods, AdaptGCD integrates learnable adapters parallel to the feed-forward layer of the ViT block and keeps the parameters from backbone frozen to preserve the pretrained knowledge. Considering the discrepancy of supervision information between the old and new classes, the adapter is further extended to a Multi-Expert Adapter structure (MEA), which encompasses multiple adapter experts and employs a routing layer to schedule the usage of these experts. In order to guide the experts to specialize in data from old or new classes, a route assignment constraint is designed. In summary, the key contributions are listed as follows:
@@
(i) Our proposal is the first work to introduce the adapter tuning into the GCD task. Such combination is elaborate, as it mechanistically allows preserving the general knowledge of pretrained backbone and meanwhile enhancing the adaptability to the downstream GCD task. We believe this could provide crucial insight for the future work of GCD.
@@
(ii) To reduce the information interference between old and new classes, we employ a multi-expert adapter structure and design a route assignment constraint to separate the data from old and new classes into different expert groups.
@@
(iii) We conduct extensive experiments on 7 popular datasets. Impressively, the adapter improves the GCD performance by a large margin, which highlights the applicability of adapter tuning on the GCD task. Furthermore, the results also confirm the effectiveness of multi-expert adapter structure with route assignment constraint.
2 Related Work
@@
Generalized Category Discovery. Generalized Category Discovery (GCD, DBLP:conf/cvpr/gcdVazeHVZ22 ; DBLP:conf/nips/VazeVZ23 ; DBLP:conf/nips/RastegarDS23 ; DBLP:conf/cvpr/promptcalgcdZhangKSNCK23 ; wang2024sptnet ; DBLP:conf/nips/BaiLWCMLZFWH23 ) aims to classify the unlabeled images across both old and new classes. The pioneering work DBLP:conf/cvpr/gcdVazeHVZ22 approaches the goal via refining the representation from the pretrained model with contrastive learning and assigning the classification results by semi-supervised $k$-means clustering. Some concurrent works DBLP:conf/cvpr/PuZS23 ; DBLP:conf/iccv/gmmgcdZhaoW023 ; DBLP:conf/cvpr/promptcalgcdZhangKSNCK23 also follow the contrastive representation learning paradigm. Distinguished from these methods, the parametric approach SimGCD DBLP:conf/iccv/simgcdWenZQ23 introduces a learnable classifier and self-distillation to supervise the classifier, achieving boosting performance. Despite the excellent performance, most of these works DBLP:conf/cvpr/gcdVazeHVZ22 ; DBLP:conf/iccv/simgcdWenZQ23 adopt the simple fine-tuning strategy only updating the last few blocks of the vision transformer, which disrupts the pretrained knowledge. Hence, this work is devoted to seeking an effective fine-tuning strategy for the GCD task to preserve the priority and enhance adaptability to the downstream tasks simultaneously.
@@
Fine-Tuning Strategy. The mainstream fine-tuning approaches includes partial tuning, prompt tuning and adapter tuning methods. Partial tuning methods freeze most of the backbone and finetune a small portion of parameters, like linear heads DBLP:conf/cvpr/IofinovaPKA22 , MLP heads DBLP:journals/corr/abs-2003-04297 , or several blocks of the backbone DBLP:conf/cvpr/HeCXLDG22 . Although this strategy is widely used in GCD, it still falls short of maintaining the network’s generalization capabilities. Prompt tuning DBLP:conf/eccv/vptJiaTCCBHL22 ; DBLP:conf/iccv/E2VPTHanWCCWQL23 prepends a set of learnable vectors to the input and only updates these prompts during fine-tuning. Despite their admirable performance, they present limited adaptability when a large gap exists between the downstream task and the upstream pretraining task DBLP:conf/eccv/vptJiaTCCBHL22 . Adapter tuning DBLP:conf/nips/adaptformerChenGTWSWL22 ; DBLP:journals/corr/abs-2302-08106 integrates shallow modules with the fixed backbone to protect the priority in pretrained models. Recent works DBLP:conf/acl/Ye020 ; DBLP:conf/cvpr/LiLB22 ; DBLP:conf/aaai/XinDWLY24 also show its flexibility to adapt to different downstream tasks. Considering the excellent mechanism of adapter tuning, this work further concentrates on the application and modification of the adapter-tuning-based strategies in GCD.
@@
Mixture of Expert. The Mixture of Experts (MoE) is a machine learning technique employing route layer to segment a single task space into multiple subtasks DBLP:journals/corr/abs-2402-08562 ; DBLP:journals/corr/abs-2311-03285 ; DBLP:journals/corr/abs-2310-18339 ; DBLP:journals/corr/abs-2310-18547 ; DBLP:journals/corr/abs-2307-13269 . Each subtask is specifically managed by individual expert networks and the prediction result is aggregation of the solutions provided by these experts. In recent years, there appear some studies focusing on the fusion of the MoE with Adapters or LoRA DBLP:journals/corr/abs-2310-18339 ; dou2024loramoe . For instance, LoRAMoE dou2024loramoe introduces a MoE-style plugin and a localized balancing constraint to segregate world knowledge and downstream information into two distinct sets of experts. Inspired by LoRAMoE, the proposed approach also introduces multi-expert technology to handle data from new and old classes in diverse experts, thus reducing interference during the learning process.
3 Method
In this work, a novel AdaptGCD framework (as depicted in Figure 1) is proposed for tackling the generalized category discovery task. In the following, we would elaborate on our proposal.
3.1 Setting of Generalized Category Discovery
@@
Assume that the training dataset is composed of the labeled and unlabeled parts defined as $\mathcal{D}^{l}=\{(\textbf{x}^{l}_{i},y^{l}_{i})\}\in\mathcal{X}\times\mathcal{Y}^{l}$ and $\mathcal{D}^{u}=\{(\textbf{x}^{u}_{i},y^{u}_{i})\}\in\mathcal{X}\times\mathcal{Y}^{u}$, where $\mathcal{Y}^{l},\mathcal{Y}^{u}$ are the label spaces of the labeled and unlabeled samples. In the GCD setting, $\mathcal{Y}^{l}\subset\mathcal{Y}^{u}$. Given $\mathcal{Y}^{n}=\mathcal{Y}^{u}\backslash\mathcal{Y}^{l}$, the categories in the $\mathcal{Y}^{l}$ are the “old” classes while those in $\mathcal{Y}^{n}$ are the “new” classes. The goal of GCD is to learn a model to categorize the samples from $\mathcal{D}^{u}$ in both the old and new classes.
3.2 Baselines: SimGCD
@@
As a simple yet effective parametric approach, SimGCD DBLP:conf/iccv/simgcdWenZQ23 is adopted as the baseline of the proposed method. It incorporates two vital losses, including representation learning loss $\mathcal{L}_{\rm rep}$ and classification objective $\mathcal{L}_{\rm cls}$.
@@
Parametric Classifier. SimGCD employs the parametric classifier in conjunction with the self-distillation paradigm for the learning of the classifier. Given the set of prototypes $\mathcal{C}=\{\textbf{c}_{1},...\textbf{c}_{K}\}$, where each one corresponds to a class and $K=|\mathcal{Y}_{u}|$ is the total number of categories. The prediction is determined by the cosine distance between the hidden features $\textbf{h}_{i}=f(\textbf{x}_{i})$ and the prototypes, that is ${p}_{i,k}=\frac{\exp\left(\frac{1}{\tau_{s}}(\textbf{h}_{i}/||\textbf{h}_{i}||_{2})^{\top}(\textbf{c}_{k}/||\textbf{c}_{k}||_{2})\right)}{\sum_{k^{\prime}}\exp\left(\frac{1}{\tau_{s}}(\textbf{h}_{i}/||\textbf{h}_{i}||_{2})^{\top}(\textbf{c}_{k^{\prime}}/||\textbf{c}_{k^{\prime}}||_{2})\right)}$, where $\tau_{s}$ indicates the temperature.
SimGCD Loss. For the representation loss, it integrates supervised contrastive learning DBLP:journals/corr/abs-2004-11362 for labeled samples with self-supervised contrastive learning DBLP:conf/icml/ChenK0H20 for all samples. Considering the classification loss, it is defined as the cross-entropy loss between the predictions and pseudo-labels or ground-truth labels. All in all, the objective for SimGCD $\mathcal{L}_{\rm gcd}$ is described as the sum of representation loss $\mathcal{L}_{rep}$ and classification loss $\mathcal{L}_{cls}$, i.e., $\mathcal{L}_{\rm gcd}=\mathcal{L}_{rep}+\mathcal{L}_{cls}$.
@@
Partial Tuning. The effectiveness of the SimGCD framework largely stems from pretrained model. To take advantage of the generic representation from upstream tasks (like DINO DBLP:conf/iccv/CaronTMJMBJ21 ), the SimGCD framework only unfreezes the pretrained parameters of the last block for fine-tuning. However, as mentioned previously, this method might still result in information loss from the pretrained backbone, thereby reducing the generality of features.
3.3 Multi-Expert Adapter

### v2

@@ Abstract
Different from the traditional semi-supervised learning paradigm that is constrained by the close-world assumption, Generalized Category Discovery (GCD) presumes that the unlabeled dataset contains new categories not appearing in the labeled set, and aims to not only classify old categories but also discover new categories in the unlabeled data. Existing studies on GCD typically devote to transferring the general knowledge from the self-supervised pretrained model to the target GCD task via some fine-tuning strategies, such as partial tuning and prompt learning. Nevertheless, these fine-tuning methods fail to make a sound balance between the generalization capacity of pretrained backbone and the adaptability to the GCD task. To fill this gap, in this paper, we propose a novel adapter-tuning-based method named AdaptGCD, which is the first work to introduce the adapter tuning into the GCD task and provides some key insights expected to enlighten future research. Furthermore, considering the discrepancy of supervision information between the old and new classes, a multi-expert adapter structure equipped with a route assignment constraint is elaborately devised, such that the data from old and new classes are separated into different expert groups. Extensive experiments are conducted on 7 widely-used datasets. The remarkable performance improvements highlight the efficacy of our proposal and it can be also combined with other advanced methods like SPTNet for further enhancement.
Index Terms:
Generalized Category Discovery, Semi-supervised Learning, Open-World Learning
I Introduction
@@
Recent advancements in deep learning have been largely driven by the availability of high-quality labeled data, which incurs significant costs associated with manual annotation. To tackle this challenge, Semi-Supervised Learning (SSL) [1, 2, 3, 4, 5] has gained considerable attention as it effectively combines a limited number of labeled samples with extensive unlabeled data for model training. While recent semi-supervised studies [6, 7, 8, 9, 10, 11] show a substantial progression, they are typically constrained by the closed-world assumption, i.e., labels are provided for each category that the model needs to classify [12]. The constraint hinders the scalability of these methods. To relax this assumption, Generalized Category Discovery (GCD) [13] presumes that unlabeled dataset contains new categories not appearing in the labeled set, as shown in Fig. 1. The goal of GCD is not only to classify the old categories but also to discover the new categories in the unlabeled data. As a promising research topic with the potential to broaden the application scope of SSL, GCD has garnered significant attention from the community in recent years.
Figure 1: The description of the generalized category discovery (GCD) task. In this context, the unlabeled dataset contains new classes that are not present in the labeled set.
@@
The primary challenge in GCD is learning the expressive representation applicable to both old and new classes [13]. Due to the scarcity of labels in GCD tasks, the model trained from scratch is prone to overfit the labeled data in old classes, consequently disrupting the generality of representation [14]. As an alternative, the existing studies in GCD broadly fine-tune the self-supervised pretrained models, like representation learning model DINO [15], to transfer the strong general knowledge from the pretrained backbone to the target dataset in GCD task. In this process, a concern that is key yet often overlooked is to strike a balance between the generalization capacity from the upstream knowledge and the adaptability to the downstream GCD task, which is largely influenced by the fine-tuning strategy [16, 17, 18, 19].
@@
Up to now, two fine-tuning strategies are widely utilized in GCD: the partial tuning and the prompt tuning. The straightforward partial tuning strategy only updates the parameters from the last few blocks while the rest parameters in the backbone are frozen [13]. Such approaches potentially impair the generalization ability with limited annotations since the pretrained knowledge embodied in the backbone is disrupted. Conversely, the prompt tuning methods [18, 20] like PromptCAL [21] and SPTNet [14] freeze all network parameters in the backbone and incorporate a learnable prompt at the input for training. Although effectively preserving the knowledge of the pretrained network, unfortunately, it is demonstrated that the prompt-based approaches are not robust to the bad data quality [14], and at the same time present limited adaptability when the downstream task differs from the upstream pretraining task [18]. This is particularly evident during the adaptation process from representation learning methods to the downstream classification task, resulting in the performance on GCD tasks far from satisfactory.
@@
Luckily, as one of the most characteristic fine-tuning approaches, adapter tuning [22, 23, 24, 25, 26] may provide a promising solution to the issues mentioned above. The adapter tuning leaves the parameters in pretrained model fixed and brings in a modest number of task-specific parameters via bottleneck adapter modules [27]. Obviously, on mechanism, the adapter naturally fits for the GCD task. It could preserve pretrained knowledge nondestructively by maintaining the parameters in the backbone [28, 29, 30]. Additionally, previous work has demonstrated its capability to adapt from upstream feature learning tasks to various downstream tasks [22, 31, 32, 33, 34]. Surprisingly, the adapter has not yet attracted much attention on GCD task. When introducing adapter into GCD, an additional concern arises accordingly, i.e., the information imbalance between old and new classes within adapter [12]. Due to the existence of supervised loss for labeled data, the adapter receives more supervision from old classes than the new ones, thereby resulting in a bias towards old classes. Under this framework, how to render old classes and new classes non-interfering in GCD task deserves further exploration.
@@
In this study, we propose a novel multi-expert adapter tuning method AdaptGCD for generalized category discovery. Unlike the previous partial tuning or prompt-like methods, AdaptGCD integrates learnable adapters parallel to the feed-forward layer of the ViT block and keeps the parameters from the backbone frozen to preserve the pretrained knowledge. Considering the discrepancy of supervision information between the old and new classes, the adapter is further extended to a Multi-Expert Adapter structure (MEA), which encompasses multiple adapter experts and employs a routing layer to schedule the usage of these experts. In order to guide the experts to specialize in data from the old or new classes, a route assignment constraint is further designed. It is also noted that the proposal could act as a plugin, allowing integration with other cutting-edge methods like SPTNet, to achieve the superior performance. In summary, the key contributions are listed as follows:
@@
  • Our proposal is the first work to introduce the adapter tuning into the GCD task. Such combination is elaborate, as it mechanistically allows preserving the general knowledge of pretrained backbone and meanwhile enhancing the adaptability to the downstream GCD task. It can also serve as a plug-and-play module to further improve the performance of the advanced method. We believe this could provide crucial insight for the future work of GCD.
@@
  • To reduce the information interference between old and new classes, we employ a multi-expert adapter structure and design a route assignment constraint to separate the data from old and new classes into different adapter expert groups, alleviating the bias towards old classes.
@@
  • We conduct extensive experiments on 7 popular datasets. Impressively, the adapter improves the GCD performance by a large margin, which highlights the applicability of adapter tuning on the GCD task. Furthermore, the results also confirm the effectiveness of multi-expert adapter structure with route assignment constraint.
The rest of the paper is organized as follows. First, in Section II, we provide a brief review of related work. Next, Section III introduces the widely used preliminary in the GCD task. In Section IV, we elaborate on the details of the proposed AdaptGCD framework. Subsequently, extensive experiments are conducted, and additional experimental analyses are presented in Section V. Finally, we summarize the paper in Section VI.
II Related Work
II-A Semi-Supervised Learning
To date, advancements in deep learning have largely depended on high-quality labeled data, which incurs substantial costs due to manual annotation. To reduce this reliance, researchers have turned to semi-supervised learning (SSL) [35, 36, 37, 38, 39, 40, 41, 42, 43, 1, 2, 3, 4, 5], which leverages a limited number of labeled samples alongside extensive unlabeled datasets for model training. Modern SSL algorithms primarily consist of two paradigms: consistency regularization [44, 45] and pseudo-labeling methods [46, 47]. Consistency regularization requires the model to learn consistent representations or predictions of the same image across various augmentations or training phases, with notable works including temporal ensembling [44] and mean teacher [45]. In contrast, pseudo-labeling methods [46] estimate reliable pseudo-labels for unlabeled data. Recently, FixMatch [6] has emerged as a combination of these paradigms, estimating pseudo-labels from weakly augmented images and minimizing cross-entropy loss between these pseudo-labels and predictions from strongly augmented images. This approach has garnered attention, leading to subsequent works like MixMatch [7] and FlexMatch [8], which enhance the scalability and performance. Additionally, another promising direction is to combine contrastive learning [48] with semi-supervised learning to enhance model generalization, thereby making better use of unlabeled data. Despite progress in SSL, the mainstream SSL methods still maintain a closed-world assumption, presuming that labeled instances are available for all possible categories in the unlabeled dataset, which limits their application in more realistic scenarios.
II-B Generalized Category Discovery
@@
As an extension of semi-supervised learning in open scenarios, Generalized Category Discovery (GCD, [13, 49, 50, 21, 14, 51, 52]), also known as open-world semi-supervised learning, does not assume that the unlabeled dataset comes from the same class set as the labeled dataset, posing a greater challenge for designing an effective model to classify unlabeled images across both old and new classes. The pioneering work [13] achieves this goal by refining representations from the pretrained model using contrastive learning and assigning classification results through semi-supervised $k$-means clustering. Some concurrent works [53, 54, 21] also follow the contrastive representation learning paradigm. Distinct from these methods, the parametric approach SimGCD [55] introduces a learnable classifier and self-distillation to supervise the classifier, achieving boosting performance. Furthermore, subsequent works have made further improvements to solve the issues of forgetting during training [52] and the desynchronization of teacher-student network knowledge [56]. Despite excellent performance, most of these studies [13, 55] adopt simple fine-tuning strategies that only update the last block of the vision transformer, which disrupts pretrained knowledge. Recently emerged prompt-based methods, such as PromptCAL [21] and SPTNet [14], can alleviate the disruption of backbone to some extent, but they still lack adaptability to various tasks and datasets. Hence, this work is devoted to seeking an effective fine-tuning strategy for the GCD task to preserve priority and enhance adaptability to downstream tasks simultaneously.
II-C Fine-Tuning Strategy
@@
The mainstream fine-tuning approaches include partial tuning, prompt tuning and adapter tuning methods. Partial tuning methods freeze most of the backbone and fine-tune a small portion of parameters, like linear heads [57], MLP heads [58], or several blocks of the backbone [59]. Although this strategy is widely used in GCD, it still falls short of maintaining the network’s generalization capabilities. Prompt tuning [18, 60] prepends a set of learnable vectors to the input and only updates these prompts during fine-tuning. Despite their admirable performance, these methods present limited adaptability when a large gap exists between the downstream task and the upstream pretraining task [18], especially during the adaptation from upstream representation learning tasks to downstream classification tasks. Adapter tuning, including AdaptFormer [22], RepAdapter [31], etc., integrates shallow modules with the fixed backbone to protect the priority in pretrained models. Recent works [32, 33, 34] also show its flexibility to adapt to different downstream tasks. Considering the excellent mechanism of adapter tuning, this work further concentrates on the application and modification of the adapter-tuning-based strategies in GCD.
II-D Mixture of Expert
@@
The Mixture of Experts (MoE) is a machine learning technique employing route layer to segment a single task space into multiple subtasks [61, 62, 63, 64, 65]. Each subtask is specifically managed by individual expert networks and the prediction result is aggregation of the solutions provided by these experts. In recent years, there appear some studies focusing on the fusion of the MoE with Adapters or LoRA [63, 66]. For instance, LoRAMoE [66] introduces a MoE-style plugin and a localized balancing constraint to segregate world knowledge and downstream information into two distinct sets of experts. Inspired by LoRAMoE, the proposed approach also introduces multi-expert technology to handle data from new and old classes in diverse experts, thus reducing interference during the learning process.
III Preliminary
In this section, we will introduce the basic setting for the generalized category discovery task and then elaborate on the pipeline of the baseline SimGCD.
III-A Setting of Generalized Category Discovery
@@
Assume that the training dataset is composed of the labeled and unlabeled parts defined as $\mathcal{D}^{l}=\{(\mathbf{x}^{l}_{i},y^{l}_{i})\}\in\mathcal{X}\times\mathcal{Y}^{l}$ and $\mathcal{D}^{u}=\{(\textbf{x}^{u}_{i},y^{u}_{i})\}\in\mathcal{X}\times\mathcal{Y}^{u}$, where $\mathcal{Y}^{l},\mathcal{Y}^{u}$ are the label spaces of the labeled and unlabeled samples. $K=|\mathcal{Y}^{u}|$ is the total number of categories. In GCD setting, $\mathcal{Y}^{l}\subset\mathcal{Y}^{u}$. Given $\mathcal{Y}^{n}=\mathcal{Y}^{u}\backslash\mathcal{Y}^{l}$, the categories in the $\mathcal{Y}^{l}$ are denoted as the “old” classes while those in $\mathcal{Y}^{n}$ are the “new” classes. The ultimate goal of the generalized category discovery task is to learn a model to categorize the samples from $\mathcal{D}^{u}$ in both the old and new classes.
III-B Baseline: SimGCD
@@
As a simple yet effective parametric approach in the field of GCD, SimGCD [55] is adopted as the baseline of the proposed method. It incorporates two vital losses, including the representation learning loss $\mathcal{L}_{\rm rep}$ and the classification objective $\mathcal{L}_{\rm cls}$.
Figure 2: The framework of our proposed AdaptGCD. It contains two critical modules: the multi-expert adapter (MEA) structure and the route assignment constraint. The MEA introduces multiple adapter experts and the route assignment constraint, including the balanced assignment loss $\mathcal{L}_{ba}$ and category-aware balanced assignment loss $\mathcal{L}_{cba}$, guides the allocation of these experts.
1) Representation Learning Loss: The representation learning loss integrates supervised contrastive learning [67] for labeled samples with self-supervised contrastive learning [48] for all samples. Given two random augmented views of the same image $\textbf{x}_{i}$ and $\textbf{x}^{\prime}_{i}$ in a mini-batch $\mathcal{B}$, the self-supervised contrastive loss $\mathcal{L}_{\rm rep}^{u}$ and supervised contrastive loss $\mathcal{L}_{\rm rep}^{l}$ are represented as Eq. (1).
$\mathcal{L}_{\rm rep}^{u}=$ $\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}-\log\frac{e^{\textbf{z}^{\top}_{i}\textbf{z}^{\prime}_{i}/\tau_{u}}}{\sum_{i}^{i\neq n}{e^{\textbf{z}^{\top}_{i}\textbf{z}^{\prime}_{n}/\tau_{u}}}},$ (1)
$\mathcal{L}_{\rm rep}^{s}=$ $\frac{1}{|\mathcal{B}^{l}|}\sum_{i\in\mathcal{B}^{l}}\sum_{q\in|\mathcal{N}_{i}|}-\log\frac{e^{\textbf{z}^{\top}_{i}\textbf{z}^{\prime}_{q}/\tau_{c}}}{\sum_{i}^{i\neq n}{e^{\textbf{z}^{\top}_{i}\textbf{z}^{\prime}_{n}/\tau_{c}}}},$
where $\textbf{z}_{i}=g(f(\textbf{x}_{i}))$ represents $\ell_{2}$-normalised feature, $f(\cdot)$, $g(\cdot)$ denote the backbone and projection head, respectively. The hyper-parameter $\tau_{u},\tau_{c}$ symbolize temperature values, $\mathcal{N}_{i}$ indicates all other labeled data with the same label as $\textbf{x}_{i}$ in the same batch, and $\mathcal{B}_{l}$ corresponds to the labeled subset of $\mathcal{B}$. Subsequently, the total loss in representation learning merges the two components with a balancing factor $\lambda$, denoted by $\mathcal{L}_{rep}=(1-\lambda)\mathcal{L}_{rep}^{u}+\lambda\mathcal{L}_{rep}^{s}$.
@@
2) Classification Objective: SimGCD employs the parametric classifier in conjunction with the self-distillation paradigm for the learning of the classifier. Given the set of category prototypes $\mathcal{C}=\{\textbf{c}_{1},...\textbf{c}_{K}\}$, where each one corresponds to a class and $K=|\mathcal{Y}^{u}|$ is the total number of categories. As depicted in Eq. (2), the prediction is determined by the cosine distance between the hidden features $\textbf{h}_{i}=f(\textbf{x}_{i})$ and the class prototypes.
$\textbf{p}_{i}^{(k)}=\frac{\exp\left(\frac{1}{\tau_{s}}(\textbf{h}_{i}/||\textbf{h}_{i}||_{2})^{\top}(\textbf{c}_{k}/||\textbf{c}_{k}||_{2})\right)}{\sum_{k^{\prime}}\exp\left(\frac{1}{\tau_{s}}(\textbf{h}_{i}/||\textbf{h}_{i}||_{2})^{\top}(\textbf{c}_{k^{\prime}}/||\textbf{c}_{k^{\prime}}||_{2})\right)}.$ (2)
In a similar manner, the soft pseudo-label $\textbf{q}^{\prime}_{i}$ is generated using another view $\textbf{x}^{\prime}_{i}$ with a sharper temperature $\tau_{r}$. Subsequently, the classification losses are defined as the cross-entropy loss between the predictions and pseudo-labels or ground-truth labels. The formulation is represented in Eq. (3).
$\mathcal{L}_{\rm cls}^{u}=$ $\frac{1}{|\mathcal{B}|}\sum_{i\in|\mathcal{B}|}\ell(\textbf{q}^{\prime}_{i},\textbf{p}_{i})-\epsilon H(\overline{\textbf{p}}),$ (3)
$\mathcal{L}_{cls}^{s}=$ $\frac{1}{|\mathcal{B}^{l}|}\sum_{i\in\mathcal{B}^{l}}\ell(\sigma({y}_{i}),\textbf{p}_{i}),$
where $\ell(\cdot,\cdot)$ refers to the cross entropy loss, $\sigma(\cdot)$ denote the one-hot operation for labels, entropy regularizer $H(\bar{\textbf{p}})=-\sum_{k}\bar{\textbf{p}}^{(k)}\log\bar{\textbf{p}}^{(k)}$ and $\bar{p}=\frac{1}{2|\mathcal{B}|}\sum_{i\in\mathcal{B}}(\textbf{p}_{i}+\textbf{p}^{\prime}_{i})$. These elements are integrated into the classification objective as $\mathcal{L}_{cls}=(1-\lambda)\mathcal{L}_{\rm cls}^{u}+\lambda\mathcal{L}_{\rm cls}^{s}$. All in all, the objective for SimGCD is described as $\mathcal{L}_{sgcd}=\mathcal{L}_{\rm rep}+\mathcal{L}_{\rm cls}$.
@@
3) Partial Tuning: The effectiveness of the SimGCD framework largely stems from pretrained model. To take advantage of the generic representation from upstream tasks (like DINO [15]), the SimGCD framework only unfreezes the pretrained parameters of the last block for fine-tuning. However, as mentioned previously, this method might still result in information loss from the pretrained backbone, thereby reducing the generality of features.
IV The Proposed Method
In this work, a novel AdaptGCD framework (as depicted in Fig. 2) is proposed for tackling the generalized category discovery task. It contains two critical modules, including the multi-expert adapter structure and the route assignment constraint. In the following contents, we would elaborate on our proposed method.
IV-A Multi-Expert Adapter